CLISP(1) CLISP(1) NAME clisp−CommonLisplanguageinterpreterandcompiler SYNOPSIS clisp[-h][-mmemsize][-Mmemfile][-Llanguage][-q][-I][-iinitfile...][-c[-l] lispfile...][-xexpression] DESCRIPTION Invokesthecommonlispinterpreterandcompiler.Invokedwithoutarguments,executesareadeval-printloop,inwhichexpressionsareinturnreadfromstandardinput,evaluatedbythelisp interpreter,andtheirresultsoutputtostandardoutput.Invokedwith−c,thespecifiedlispfiles arecompiledtoabytecodethatcanbeexecutedmoreefficiently. OPTIONS -hDisplaysahelpmessageonhowtouseclisp. -mmemsize Setstheamountofmemoryclisptriestograbonstartup.Theamountmaybegivenas nnnnnnn(measuredinbytes),nnnnKornnnnKB(measuredinkilobytes)ornMor nMB(measuredinmegabytes).Defaultis2megabytes.Theargumentisconstrained between100KBand16MB.--Thisversionofclispallocatesmemorydynamically. memsizeisessentiallyignored. -Mmemfile Specifiestheinitialmemoryimage.Thismustbeamemorydumpproducedbythe saveinitmemfunction. -Llanguage Specifiesthelanguageclispusestocommunicatewiththeuser.Thismaybeenglish, deutsch,francais. -qQuiet:clispdisplaysnobanneratstartupandnogood-byemessagewhenquitting. -IILISPfriendly:clispinteractsinawaythatILISP(apopularEmacsLISPinterface)can dealwith.Currentlytheonlyeffectofthisisthatunnecessarypromptsarenotsuppressed. -iinitfile... Specifiesinitializationfilestobeloadedatstartup.Theseshouldbelispfiles(sourceor compiled). -clispfile... Compilesthespecifiedlispfilestobytecode.Thecompiledfilescanthenbeloadedinstead ofthesourcestogainefficiency. -lAbytecodelistingofthefilesbeingcompiledwillbeproduced.Usefulonlyfordebugging purposes. -xexpressions Executesaseriesofarbitraryexpressionsinsteadofaread-eval-printloop.Thevaluesof theexpressionswillbeoutputtostandardoutput.Duetotheargumentprocessingdone bytheshell,theexpressionsmustbeenclosedinsinglequotes,anddoublequotesand backslashesmustbeprecededbybackslashes. @optionfile substitutesthecontentsofoptionfileasarguments.Eachlineofoptionfileistreatedasa separateargumenttoclisp. REFERENCE Thelanguageimplementedconformsto GuyL.SteeleJr.:CommonLisp-TheLanguage. DigitalPress.1stedition1984,465pages. ("CLtL1"forshort) 19June1994 1 CLISP(1) CLISP(1) andtotheolderpartsof GuyL.SteeleJr.:CommonLisp-TheLanguage. DigitalPress.2ndedition1990,1032pages. ("CLtL2"forshort) USE helptogetsomeon-linehelp. (aproposname) liststhesymbolsrelatingtoname. (exit)or(quit)or(bye) toquitclisp. EOF(Ctrl-Z) toleavethecurrentread-eval-printloop. arrowkeys foreditingandviewingtheinputhistory. Tabkey tocompletethesymbol’snameyouarejusttyping. FILES lisp.exe mainexecutable lispinit.mem initialmemoryimage config.lsp site-dependentconfiguration *.lsplispsource *.faslispcode,compiledbyclisp *.liblispsourcelibraryinformation,generatedandusedbytheclispcompiler ENVIRONMENT CLISPLANGUAGE specifiesthelanguageclispusestocommunicatewiththeuser.Thevaluemaybe english,deutsch,francaisanddefaultstoenglish.The-Loptioncanbeusedtoover-ridethisenvironmentvariable. COMSPEC(DOS,OS/2implementationsonly) isusedtofindthecommandinterpretercalledbythefunctionshell. TERM specifiestheterminalemulationclisprelieson.IfyouhaveANSI.SYSloaded,possible valuesareansi,ansi-color-2,ansi-color-3andmono. TERMCAP shouldbesettotheslashifiedfilenameoftheterminalcapabilitiesdatabasetermcap.dat. SEEALSO cmulisp(1),emacs(1). BUGS Thefunctioninspectisnotimplemented. OnlyveryfewextensionsfromCLtL2aresupported. Noon-linedocumentationbeyondaproposanddescribeisavailable. Stackoverflowabortstheprogramungracefully,witharegisterdump. PressingControl-Cmaynotinterruptclispineverysituation. 19June1994 2 CLISP(1) CLISP(1) CallingthefunctionEXECUTEonbatchfilescrashesthemachine. PROJECTS Writingon-linedocumentation. Buildingaforeignfunctioninterface(abilitytocallCcodedirectly). Writeinspect. Enhancethecompilersuchthatitcaninlinelocalfunctions. Specifyaportablesetofwindowandgraphicsoperations. AUTHORS BrunoHaible<haible@ma2s2.mathematik.uni-karlsruhe.de>andMichaelStoll. 19June1994 3